from a base elevation of 7300 ft, a mountain peak in colorado rises to a summit elevation of 14807 ft over a horizontal distance of 15840 ft find the grade of the peak
Answer by bucky(2189)

In this problem the grade (call it G) is the slope and the slope is defined as the vertical rise divided
by the horizontal run.
.
So for this problem, the vertical rise (or elevation) is 14807 feet ... and the horizontal run
is 15840. Note that both these dimensions are in the common units of feet. The units of the
denominator and the numerator must be the same.
.
So for this problem you have:
.
G = (elevation)/(horizontal distance) = 14807/15840 = 0.934785
.
or the Grade is approximately 0.935 ... meaning a 0.935 foot rise for every 1 foot
you travel horizontally.
.
Hope this helps you to understand the problem and how to work it.
.